OpenStudy (loser66):

Prove that the map \[Z_2\rightarrow Z_2\\[r]\rightarrow [r]^2\] for \(r\in Z\), is a ring homomorphism Please, help

OpenStudy (anonymous):

Now what? hahah

terenzreignz (terenzreignz):

Hmm... it's not showing up. I guess you mean \[\Large \left.\begin{matrix} \mathbb{Z}_2 &\rightarrow & \mathbb{Z}_2\\ r & \mapsto & r^2\end{matrix}\right.\]

terenzreignz (terenzreignz):

Is that it?

OpenStudy (loser66):

[r] not just r

OpenStudy (anonymous):

hahah

OpenStudy (loser66):

congruence class

terenzreignz (terenzreignz):

Oh...

terenzreignz (terenzreignz):

Let me see if I can learn what in the blazes a congruence class in within a minute. That is, if you don't educate me first, LOSER :P (haha... peace ^^)

OpenStudy (loser66):

\[Z_2 =\{[0],[1]\}\]

terenzreignz (terenzreignz):

So what is a bloody congruence class, then? :D

terenzreignz (terenzreignz):

\[\Large \left.\begin{matrix} \mathbb{Z}_2 &\rightarrow & \mathbb{Z}_2\\ [r] & \mapsto & [r]^2\end{matrix}\right.\]

OpenStudy (loser66):

in Z_2, [r]^2= Z_2 itself

terenzreignz (terenzreignz):

Yeah, but it would help if you tell me what [r] really means, you know, its mathematical definition?

OpenStudy (loser66):

[r]= {km +a | k, m in Z, 0<= a<m}

OpenStudy (loser66):

and a = r (mod m)

OpenStudy (loser66):

for example in mod 2, [1]= {1,3,5,......}

OpenStudy (loser66):

because 2|3 remain 1 like 2|5 and so on so that [r] = {1,3,5,......}

OpenStudy (anonymous):

What's this?

terenzreignz (terenzreignz):

Okay, so the gist of it is, in mod 2, which is in Z_2, [1] is all the numbers that give a remainder of 1 when divided by 2?

OpenStudy (loser66):

yes

OpenStudy (loser66):

@EngrChong that is group theory

OpenStudy (anonymous):

I dont understand anything hahaha

terenzreignz (terenzreignz):

Okay.... and what does [r]^2 mean?

OpenStudy (loser66):

mean {1^2,3^2,5^2......} but in mod 2, it is Z_2

OpenStudy (loser66):

because 3^2=9 and 2|9 remain 1 also

terenzreignz (terenzreignz):

So [r] and [r]^2 are sets, yes?

OpenStudy (loser66):

[r]^2 = [r][r]=[r^2]

OpenStudy (loser66):

yes,

terenzreignz (terenzreignz):

And how do you define [r] + [s] for instance? And [r][s] ?

terenzreignz (terenzreignz):

See, the way I was taught, Z_2 is simply the set {0,1}

OpenStudy (loser66):

[r]+[s]=[r+s]

OpenStudy (loser66):

but it is not equal r +s, right?

OpenStudy (loser66):

for example: [1] + [0]= [1+0]=[1] , but [1]\(\neq \)1 , [1] is a set of numbers whose remainder is 1 in mod2, 1 is just one of its element.

OpenStudy (loser66):

Oh, I see, since you study group theory while my course is abstract algebra.

terenzreignz (terenzreignz):

So the members of Z_2 are sets?

OpenStudy (loser66):

yes

terenzreignz (terenzreignz):

That's weird

terenzreignz (terenzreignz):

Let's see... Let [r] and [s] be elements of Z_2

terenzreignz (terenzreignz):

f([r] + [s]) = f([r+s]) = [r+s]^2

terenzreignz (terenzreignz):

(I still don't know what I'm doing with this, mind you)

terenzreignz (terenzreignz):

Now, \[\Large [r+s]^2 = \left\{x^2 \quad | \quad x\equiv(r+s)(\mod 2)\right\}\] right?

OpenStudy (loser66):

yes

OpenStudy (loser66):

for you to see what the definition is.

OpenStudy (loser66):

they always define Z_n elements are [ ] congruence class of remainders, not just remainders

terenzreignz (terenzreignz):

Wait... backtrack...\[\Large [r+s]^2 =[(r+s)^2]= \left\{x \quad | \quad x\equiv(r+s)^2(\mod 2)\right\}\]

terenzreignz (terenzreignz):

I think this is better

OpenStudy (loser66):

another proof :) at my mark

OpenStudy (loser66):

ok, next?

terenzreignz (terenzreignz):

Okay, I can sort of see where this is going.... now, we'll examine \[\Large [r]^2 + [s]^2 = [r^2 + s^2]\]

terenzreignz (terenzreignz):

\[\Large [r^2+s^2]=\left\{x \quad | \quad x\equiv(r^2+s^2)(\mod 2)\right\}\]

OpenStudy (loser66):

it's true in this case only (mod 2)

terenzreignz (terenzreignz):

It's always true. I haven't put in the specific mod 2 case yet.

terenzreignz (terenzreignz):

I will in the next step. We can add any even integer to this bit, with no problem, right? \[\Large [r^2+s^2]=\left\{x \quad | \quad x\equiv(\color{blue}{r^2+s^2})(\mod 2)\right\}\]

terenzreignz (terenzreignz):

Because... a = b(mod 2) is equivalent to a = (b+2k)(mod 2), for any integer k.

OpenStudy (loser66):

ok

terenzreignz (terenzreignz):

So why don't we add 2rs? :P \[\Large [r^2+s^2]=\left\{x \quad | \quad x\equiv(r^2\color{red}{+2rs}+s^2)(\mod 2)\right\}\]

terenzreignz (terenzreignz):

Thus... \[\Large [r^2+s^2]=\left\{x \quad | \quad x\equiv(r+s)^2(\mod 2)\right\}=[r+s]^2\]

OpenStudy (loser66):

because in mod 2 , 2|2rs so 2rs\(\equiv\) 0

terenzreignz (terenzreignz):

Exactly. So we have established that \[\Large f\left([r]+[s]\right)=[r+s]^2 = [r^2+s^2]=f([r]) + f([s])\] Right? That's one bit of the ring homomorphism down haha

terenzreignz (terenzreignz):

The other part is, I think, easier.

OpenStudy (loser66):

The first one is proving it is a ring, then this stuff, right? Ok, got you, thanks a lot. You are...extremely good. hehehe.. I didn't post the problem until I saw you online. Thank you.

terenzreignz (terenzreignz):

Again, with [r] and [s] being elements of Z_2, let's have a look at [r][s] \[\Large f([r][s]) = f([rs]) = [rs]^2\]\[\Large [rs]^2 = [(rs)^2] = [r^2s^2] = [r^2][s^2]=[r]^2[s]^2 = f([r])f([s])\] Whew :P

terenzreignz (terenzreignz):

First one was proving that it was a group homomorphism. This next one proves that it is also a ring homomorphism.

OpenStudy (loser66):

:) Still have time? I need you help on other one. But it is ok if you don't have time. :)

terenzreignz (terenzreignz):

Show it to me. If I don't have time, I'll just give you the best hint that I can :)

OpenStudy (loser66):

Let P denote the multiplicative group of positive real numbers. Prove that {1} is the ONLY finite subgroup of P.

terenzreignz (terenzreignz):

multiplicative group?

OpenStudy (loser66):

group under multiplication operator

terenzreignz (terenzreignz):

example?

OpenStudy (loser66):

for example:  \[Z_8 = \{[1].[3],[5],[7]\}\] all elements 1,3,5,7 are relative prime with 8

terenzreignz (terenzreignz):

I still don't quite get what P is... let me try to see

OpenStudy (loser66):

Let me give you the lattice of subgroup, it may trig something 

terenzreignz (terenzreignz):

o.O

terenzreignz (terenzreignz):

I don't guarantee I'll be able to help with this one... o.O

OpenStudy (loser66):

It's perfectly ok, friend. I do appreciate your patience. :) I will ask my T.A tomorrow. Thanks for the help.